OpenStudy (anonymous):

Which of the following equals 5^2? 5 10 25 125

OpenStudy (anonymous):

25

OpenStudy (anonymous):

5^2 = 5 * 5 = 5 + 5 + 5 + 5 + 5 = ?

OpenStudy (anonymous):

haha its 25 lol i have no idea why i asked this hahah

OpenStudy (anonymous):

its okay lol :P